# A modified AAA algorithm for learning stable reduced-order models from data

### 28 Dec 2023 v1

@@ 3.2.2 The real-valued AAA algorithm
$\begin{split}&\displaystyle\sum_{i=1}^{\ell}\left(\frac{w_{i}^{(\ell)}}{s-j\lambda_{i}}+\frac{(w_{i}^{(\ell)})^{*}}{s+j\lambda_{i}}\right)H(s)\approx\sum_{i=1}^{\ell}\left(\frac{h_{i}^{(\ell)}w_{i}^{(\ell)}}{s-j\lambda_{i}}+\frac{(h_{i}^{(\ell)}w_{i}^{(\ell)})^{*}}{s+j\lambda_{i}}\right)\\ &\Leftrightarrow\ \ \Big{(}\sum_{i=1}^{\ell}\frac{H(s)-h_{i}^{(\ell)}}{s-j\lambda_{i}}\Big{)}w_{i}^{(\ell)}+\Big{(}\sum_{i=1}^{\ell}\frac{H(s)-(h_{i}^{(\ell)})^{*}}{s+j\lambda_{i}}\Big{)}(w_{i}^{(\ell)})^{*}:=E^{(\ell)}(s)\approx 0\quad\forall s\in\Gamma^{(\ell)}.\end{split}$ (17)
By collecting the unknowns in the vector ${\textbf{w}}^{(\ell)}=[w_{1}^{(\ell)},(w_{1}^{(\ell)})^{*},\ldots,w_{\ell}^{(\ell)},(w_{\ell}^{(\ell)})^{*}]$, condition (17) is enforced by solving the following homogeneous least-squares problem
@@
when both are are enforced in a least-squares sense; then, by observing the conditions in (19) are linear in the unknowns $\alpha_{i}^{(\ell)}=\texttt{Re}(w_{i}^{(\ell)})$ and $\beta_{i}^{(\ell)}=\texttt{Im}(w_{i}^{(\ell)})$, we can define the following vector of unknowns, ${\textbf{x}}^{(\ell)}=[\alpha_{1}^{(\ell)},\beta_{1}^{(\ell)}\ldots\alpha_{\ell}^{(\ell)},\beta_{\ell}^{(\ell)}]$, and enforce (19) by solving the following homogeneous least-squares problem
@@
where now, all the involved quantities are real-valued and the row dimension of the matrix ${\mathbb{L}}_{\textrm{real}}^{(\ell)}$ is twice that of ${\mathbb{L}}^{(\ell)}$ since the approximation conditions on real and imaginary components are considered separately. The formal definition of the real-valued “quasi-Loewner” matrix is provided by making use of the notation $\Gamma^{(\ell)}=\{\zeta_{1},\ldots,\zeta_{2V-2\ell}\}$, for all $1\leq i\leq\ell$ and $1\leq l\leq 2V-2\ell$, as:
$\left({\mathbb{L}}_{\textrm{real}}\right)_{l,2i-1}=\texttt{Re}\left(\frac{H(\zeta_{l})-h_{i}^{(\ell)}}{\zeta_{l}-j\lambda_{i}}\right)+\texttt{Re}\left(\frac{H(\zeta_{l})-(h_{i}^{(\ell)})^{*}}{\zeta_{l}+j\lambda_{i}}\right),\ \text{and}$
$\left({\mathbb{L}}_{\textrm{real}}\right)_{l,2i}=\texttt{Im}\left(\frac{H(\zeta_{l})-h_{i}^{(\ell)}}{\zeta_{l}-j\lambda_{i}}\right)-\texttt{Im}\left(\frac{H(\zeta_{l})-(h_{i}^{(\ell)})^{*}}{\zeta_{l}+j\lambda_{i}}\right).$
Algorithm 1 summarizes the above procedure in a pseudocode format.
@@ 4.3.2 The cost function of the relaxed convex problem
${\textbf{w}}_{\textrm{opt}}=\arg\min_{\left\lVert{\textbf{w}}\right\rVert_{2}=1}\|{\mathbb{L}}{\textbf{w}}\|_{2}=\arg\min_{\left\lVert{\textbf{w}}\right\rVert_{2}=1}\|\boldsymbol{\Delta}{\textbf{e}}\|_{2}=\arg\min_{\left\lVert{\textbf{w}}\right\rVert_{2}=1}{\textbf{e}}^{*}\boldsymbol{\Delta}^{\star}\boldsymbol{\Delta}{\textbf{e}}\quad\boldsymbol{\Delta}=\textrm{diag}[D(s_{1}),\ldots,D(s_{L})].$ (64)
@@ 5 Numerical results and applications
$f_{Max}=\max_{i=v,\ldots,N}\lvert\lambda_{v}/2\pi\lvert,\quad h_{Max}=\max_{v=1,\ldots,V}\lvert h_{v}\lvert,\quad h_{v}\leftarrow h_{v}/h_{Max},\quad\lambda_{v}\leftarrow\lambda_{v}/f_{Max},\quad v=1,\ldots,V.$ (68)
@@ Appendix A Appendix
A.1 Proof of Lemma 3.1
@@
A.2 Two useful state-space realizations
@@
A.2.1 A complex-valued realization with unit input map
@@
A.2.2 Conversion to real-valued realizations

### 18 Jan 2024 v2

@@ 3.2.2 The real-valued AAA algorithm
$\begin{split}&\displaystyle\sum_{i=1}^{\ell}\left(\frac{w_{i}^{(\ell)}}{s-j\lambda_{i}}+\frac{(w_{i}^{(\ell)})^{*}}{s+j\lambda_{i}}\right)H(s)\approx\sum_{i=1}^{\ell}\left(\frac{h_{i}^{(\ell)}w_{i}^{(\ell)}}{s-j\lambda_{i}}+\frac{(h_{i}^{(\ell)}w_{i}^{(\ell)})^{*}}{s+j\lambda_{i}}\right)\\ &\Leftrightarrow\ \ \Big{(}\sum_{i=1}^{\ell}\frac{H(s)-h_{i}^{(\ell)}}{s-j\lambda_{i}}w_{i}^{(\ell)}\Big{)}+\Big{(}\sum_{i=1}^{\ell}\frac{H(s)-(h_{i}^{(\ell)})^{*}}{s+j\lambda_{i}}(w_{i}^{(\ell)})^{*}\Big{)}:=E^{(\ell)}(s)\approx 0\quad\forall s\in\Gamma^{(\ell)}.\end{split}$ (17)
By collecting the unknowns in the vector ${\textbf{w}}^{(\ell)}=[w_{1}^{(\ell)},(w_{1}^{(\ell)})^{*},\ldots,w_{\ell}^{(\ell)},(w_{\ell}^{(\ell)})^{*}]^{T}$, condition (17) is enforced by solving the following homogeneous least-squares problem
@@
when both are are enforced in a least-squares sense; then, by observing the conditions in (19) are linear in the unknowns $\alpha_{i}^{(\ell)}=\texttt{Re}(w_{i}^{(\ell)})$ and $\beta_{i}^{(\ell)}=\texttt{Im}(w_{i}^{(\ell)})$, we can define the following vector of unknowns, ${\textbf{x}}^{(\ell)}=[\alpha_{1}^{(\ell)},\beta_{1}^{(\ell)}\ldots\alpha_{\ell}^{(\ell)},\beta_{\ell}^{(\ell)}]^{T}$, and enforce (19) by solving the following homogeneous least-squares problem
@@
where now, all the involved quantities are real-valued and the row dimension of the matrix ${\mathbb{L}}_{\textrm{real}}^{(\ell)}$ is twice that of ${\mathbb{L}}^{(\ell)}$ since the approximation conditions on real and imaginary components are considered separately. The formal definition of the real-valued “quasi-Loewner” matrix is explicitly provided in Appendix A.1. Algorithm 1 summarizes the above procedure in a pseudocode format.
@@ 4.3.2 The cost function of the relaxed convex problem
${\textbf{w}}_{\textrm{opt}}=\arg\min_{\left\lVert{\textbf{w}}\right\rVert_{2}=1}\|{\mathbb{L}}{\textbf{w}}\|_{2}=\arg\min_{\left\lVert{\textbf{w}}\right\rVert_{2}=1}\|\boldsymbol{\Delta}{\textbf{e}}\|_{2}=\arg\min_{\left\lVert{\textbf{w}}\right\rVert_{2}=1}{\textbf{e}}^{\star}\boldsymbol{\Delta}^{\star}\boldsymbol{\Delta}{\textbf{e}}\quad\boldsymbol{\Delta}=\textrm{diag}[D(s_{1}),\ldots,D(s_{L})].$ (64)
@@ 5 Numerical results and applications
$f_{Max}=\max_{v=1,\ldots,V}\lvert\lambda_{v}/2\pi\lvert,\quad h_{Max}=\max_{v=1,\ldots,V}\lvert h_{v}\lvert,\quad h_{v}\leftarrow h_{v}/h_{Max},\quad\lambda_{v}\leftarrow\lambda_{v}/f_{Max},\quad v=1,\ldots,V.$ (68)
@@ Appendix A Appendix
A.1 Explicit definition of ${\mathbb{L}}^{(\ell)}_{\textrm{real}}$
By making use of the notation $\Gamma^{(\ell)}=\{j\zeta_{1},\ldots,j\zeta_{V-\ell}\}$, for all $1\leq i\leq\ell$ and $1\leq l\leq V-\ell$, the elements of ${\mathbb{L}}_{\textrm{real}}$ are defined as follows:
$\left({\mathbb{L}}^{(\ell)}_{\textrm{real}}\right)_{l,2i-1}=\frac{-\texttt{Im}\left(H(j\zeta_{l})\right)+\texttt{Im}\left(h_{i}^{(\ell)}\right)}{-\zeta_{l}+\lambda_{i}}+\frac{\texttt{Im}\left(H(j\zeta_{l})\right)+\texttt{Im}\left(h_{i}^{(\ell)}\right)}{\zeta_{l}+\lambda_{i}},$
$\left({\mathbb{L}}^{(\ell)}_{\textrm{real}}\right)_{l,2i}=\frac{-\texttt{Re}\left(H(j\zeta_{l})\right)+\texttt{Re}\left(h_{i}^{(\ell)}\right)}{-\zeta_{l}+\lambda_{i}}+\frac{-\texttt{Re}\left(H(j\zeta_{l})\right)+\texttt{Re}\left(h_{i}^{(\ell)}\right)}{\zeta_{l}+\lambda_{i}},$
$\left({\mathbb{L}}^{(\ell)}_{\textrm{real}}\right)_{l+V-\ell,2i-1}=\frac{\texttt{Re}\left(H(j\zeta_{l})\right)-\texttt{Re}\left(h_{i}^{(\ell)}\right)}{-\zeta_{l}+\lambda_{i}}+\frac{-\texttt{Re}\left(H(j\zeta_{l})\right)+\texttt{Re}\left(h_{i}^{(\ell)}\right)}{\zeta_{l}+\lambda_{i}},\ \text{and}$
$\left({\mathbb{L}}^{(\ell)}_{\textrm{real}}\right)_{l+V-\ell,2i}=\frac{-\texttt{Im}\left(H(j\zeta_{l})\right)+\texttt{Im}\left(h_{i}^{(\ell)}\right)}{-\zeta_{l}+\lambda_{i}}+\frac{-\texttt{Im}\left(H(j\zeta_{l})\right)+\texttt{Im}\left(h_{i}^{(\ell)}\right)}{\zeta_{l}+\lambda_{i}}.$
A.2 Proof of Lemma 3.1
@@ A.1 Explicit definition of ${\mathbb{L}}^{(\ell)}_{\textrm{real}}$
A.3 Two useful state-space realizations
@@
A.3.1 A complex-valued realization with unit input map
@@
A.3.2 Conversion to real-valued realizations